OpenStudy (anonymous):

show work 8 = 8p + 13 - 3p

OpenStudy (anonymous):

whats the 1st step you will take?

OpenStudy (anonymous):

you have to combine like terms

OpenStudy (anonymous):

ok, do it

OpenStudy (anonymous):

i already have, my answer was 21, i just want to see if i did the work right.

OpenStudy (anonymous):

no its wrong post your work so i can see were you made the mistake

OpenStudy (anonymous):

8 = 8p + 13 - 3p 8 = 13 - 8p + 3p 8 = 13 - 11p i re did it and now im confused

OpenStudy (anonymous):

8 = 8p + 13 - 3p how did you get -8p and +3p?

OpenStudy (anonymous):

8 = 8p + 13 - 3p 8-13=8p-3p putlike terms together -5=5p divide -1=p

OpenStudy (anonymous):

@Cryxz cant you see that im teaching him

OpenStudy (anonymous):

i thought you combined like terms first?

OpenStudy (anonymous):

yes you do 8 = 8p + 13 - 3p you have +8p and -3p what would you get if you add them?

OpenStudy (anonymous):

5p? so 8 = 13 + 8p - 3p?

OpenStudy (anonymous):

yes 8 = 13 + 8p - 3p 8 = 13 + 5p now?

OpenStudy (anonymous):

8 = 13 + 5p 13 - 8= 13 - 13 + 5p 5 = 5p then simplify?

OpenStudy (anonymous):

13-8?

OpenStudy (anonymous):

subtract 13 from each side right?

OpenStudy (anonymous):

yes so is it -13 or +13?

OpenStudy (anonymous):

13 is positive?

OpenStudy (anonymous):

why positive?

OpenStudy (anonymous):

no, negative, it inherits the operation right.

OpenStudy (anonymous):

yes you are subtracting so it should be - so write the equation now

OpenStudy (anonymous):

8 = 13 + 5p -13 + 8= 13 - 13 + 5p -5 = 5p

OpenStudy (anonymous):

yes, now?

OpenStudy (anonymous):

divide by by 5p and get -1 as your answer. if thats the way its meant to be done, its no wonder im failing math this week.

OpenStudy (anonymous):

-1 is correct; just pay attention to + and - if you need other help let me know

OpenStudy (anonymous):

on the side of the equation, the positive 8, when anything is added to that side, for example the -13, as long as the 8 is positive it will be addition right?

OpenStudy (anonymous):

it will be addition if you are adding positive number for example if it was 8 = 5p - 13 you will have to add 13 to both sides